OpenStudy (breebree123me):

I will medal and fan! The position of a diving bell is −3/4 mile relative to sea level. The diving bell reached this position after dropping from sea level at a constant rate for 1 1/2 hours. What was the position of the diving bell relative to sea level 1 hour after it began its descent? Enter your answer in the box as a simplified fraction.

OpenStudy (caozeyuan):

s=-3/4, u=v=x,a=0,t=3/2 s=ut=3/2)*x so x is -1/2, you want to solve for s when t is 1, so s=-1/2*1=-1/2

OpenStudy (breebree123me):

so the answer in -1/2?

OpenStudy (caozeyuan):

s is distance, u is initial velocity, v is final velocity, a is accelaration, t is time

OpenStudy (caozeyuan):

yes

OpenStudy (breebree123me):

ok thanks